# Rosette formations as symmetry-breaking events: theory and experiment

### v1

@@ III Discussions
Rosettes are polarized, transient epithelial structures, observed as intermediate cellular structures in several biological contexts as different as morphogenesis, development, wound healing and cancer Harding et al. (2014); Moran and Mosimann (2023). From a geometrical point of view, these structure can be identified in tesselation of 2D section of cellular layers as vertex where five or more edges meet. The mechanism through which such a configuration can be reached is often proposed to be a topological rearrangement following the disappearing of a cell, i.e. a T2 transition, either due to ablation or cell apoptosis Harding et al. (2014). Indeed, theoretical and numerical investigations adopted this framework to generate rosettes and study their effects on the tissue dynamics. Using modified vertex models, it has been found that rosettes impact on tissue fluidity and organization (see e.g. the work of Yan and Bi et al. Yan and Bi (2019b)). However, rosette formation is not necessarily due to cell ablations. For instance, intercalations give rise to multiple T1 transitions leading to a 5 or more-fold vertex  Harding et al. (2014); Fletcher et al. (2014), while preserving the number of cells in the population.
@@
Thus, we sought for a more general driving principle behind the formation of rosettes. We started by noting that in a flat Euclidean space, the energetically optimal tiling for cell systems ( but also atomic crystals or soap films) is typically hexagonal. In such flat spaces, 3-fold vertices (T-junctions) are energetically optimal because at a 3-edge vertex, one can arrange the edges at $120^{\circ}$ angles (like for hexagons), balancing the tensions symmetrically. This configuration preserves both local and global spatial symmetry. Four edges junctions with angles around $90^{\circ}$ are possible, but result in unstable configurations without special conditions (i.e. peculiar disposition of the line tensions). Indeed, generalizing Plateau’s laws for soap films, at equilibrium the angles at a vertex depend on (i) the number of edges meeting, (ii) the tensions in each edge, and (iii) the local curvature of the space. Thus, higher-fold junctions could in principle become favorable for specific anisotropic tension configurations and/or where curvature affects the local angular sum at vertices. Positive curved space corresponds to sum of angles exceeding $360^{\circ}$. As the natural meeting angles between edges change, in this scenario rosettes can become stable because distributing larger angles over more edges reduces mechanical stress.
To test our intuition, we considered a population of cells described by an Active Voronoi model subject to the standard Vertex energy in Eq. 3, moving in a box with periodic boundary conditions, i.e. on a toroid. Such energy is known to produce a regular hexagonal tiling for a value of the target shape factor of $s_{0}\simeq 3.72$ and a solid to liquid-like transition for $s_{0}\simeq 3.82$. In the solid region, the system exhibits cells with both pentagonal and heptagonal shapes (see polygon distributions in Figure 1). This behavior can be understood by means of the Euler’s rule for a tessellation: $V-E+F=\chi$, where $V$ is the number of vertices, $E$ the number of edges, $F$ the number of faces, and $\chi$ is the Euler characteristic Del Bono et al. (2024) (note that for a flat plane, $\chi=0$, while for a sphere it is 2). If each vertex connects three edges, it follows that the average number of sides per face must be 6, leading to a preference for hexagonal patterns. Note that the presence of more than three edge vertices reduces the average.
Values of the shape factor different from $p_{0}^{hex}$ introduces frustrations in the system, which are compensated by the introduction of defects such as pentagons and heptagons. Each cell’s deviation from hexagonal geometry inserts a topological charge, , whose sum in flat space ($\chi=0$) must vanish. Such frustration is not sufficient to account for the presence of rosettes (see Figure 1).
Indeed, in flat surfaces, high-order vertices tend to resolve back into 3-fold vertices. In presence of a potential, rosettes can persist and balance mechanical forces and curvature constraints. Moreover, positive curvature induces cells to tighten, which reflects in cells with low number of sides (see polygon distributions in Figure 1 and  2).
@@
Introducing the external harmonic field via Eq. 4, we inserted the simplest positive correction the curvature. To accommodate such curvature, the system tends to introduce topological defects that locally reduce angular excess, forming rosette like structures (see snapshots in Figure 2). Euler’s characteristic becomes $\chi>0$, modifying the balance of defects: an excess of positive topological charge is required.
Note that from a topological point of view, the necessity of these defects in curved geometries is explained by the Gauss-Bonnet theorem, which links the integrated Gaussian curvature of a surface to the total topological defect charge. For instance, to effectively tile a sphere, one has to introduce 5-fold disclinations as is the case for physical structures like fullerenes, viral capsids, and colloidal crystals on droplets Drozdowski and Schwarz (2024).
@@
In ideal, infinite crystalline systems with flat geometry, translational symmetry ensures that the structure can be shifted in space without altering its physical properties. However, the introduction of curvature and the accompanying defects break this symmetry. On curved surfaces, there is no global translational symmetry—each point has different geometric properties. Defects such as 5- and 7-fold polygons are therefore indicators of the breaking of translational invariance. In fact, this symmetry breaking is a hallmark of systems constrained by global geometry and is essential to the emergence of localized, inhomogeneous physical properties.
@@
In the present work, we produced the breaking of the space translational invariance that generates topological defects via the introduction of either a potential energy landscape or a leader cell with different size with respect to the rest of the population. In the former case, the external potentials used can mimic the collective effects of apical actin filament networks. Indeed, actin filaments play a fundamental role in shaping the mechanical environment of cells, particularly by forming networks that exert tensile forces across the apical surface of epithelial cell layers. These supra-cellular actin networks can generate coordinated mechanical tension, contributing to tissue morphogenesis and the maintenance of epithelial integrity Fernandez-Gonzalez and Zallen (2011). In particular, it has been demonstrated how apical actomyosin cables and medial actin structures contribute to mechanical force distribution across cell layers Rauzi et al. (2010); Fernandez-Gonzalez and Zallen (2011). Similarly, the potential can simulate the effect of an actual curved surface where cells are constrained to move like in embryos and tumor spheroids Trichas et al. (2012). Note that in these real systems, there is a spontaneous breaking of the symmetry as internal cell mechanisms are responsible for processes like the accumulation of actin fibers or the shape of the embryo. Indeed, we show that the presence of a large cell in the layer suffices to insert a topological defect that provoke the formation of transient rosettes.
@@
Relying on a dynamical model, we were able to probe the stabilizing/destabilizing effects of cellular properties on the formation of rosettes. In particular, we found that cell motility and deformability influence the formation and maturation of rosettes. Indeed, our simulations highlighted an interplay between topological curvature, self propulsion velocity and target shape factor (Figure 2b-c). In presence of fast moving cells, rosettes formation needs both an increase of space curvature and a higher cell deformability. Similarly, both the size and polarization of lumen centered structures is favored in presence of a cell population in a fluid-like state. Notably, analyzing the behavior of populations of iPSC-derived neural cells growing on a 2D layer via extensive fluorescent microscopy time lapses, we confirmed our model predictions, as rosettes form in regions characterized by slow moving cells having shape factor higher than 3.9, i.e. being in a fluid like regime. Looking the motion of isolated cells, we observed that their motion is maximum when it is aligned with the direction for their elongation (Figure 6). Self-alignment interactions that produce cell displacements in the same direction along which cells are elongated is the second ingredient we identified that increases the probability of transient rosette formation (see Figure 3) and the size of lumen-centered ones (see Figure 4). In particular, it allows cells to orderly dispose in multiple layers around the lumen as shown in Figure 5.
@@
In conclusion, our work proved that rosettes are natural outcomes of geometric and mechanical responses to curvature, providing an additional framework where the interplay between defects and spatial curvature offers profound insights into how materials and surfaces are structured, with deep implications across disciplines as diverse as condensed matter physics and cosmology. To test our prediction, we measured the dynamical features of populations of iPSCs and found that neural rosettes are observed when cells populate the region of the parameter space our model predicts compatible for their formation.
@@ IV.1 Cell Culture
Human iPSCs from the WT I line Lenzi et al. (2015) were stably transfected with a construct encoding an RFP-tagged version of the nuclear FUS protein De Santis et al. (2019). iPSCs were induced to differentiate along the neuroectodermal lineage to form neural rosettes as follows. 10.000 cells/cm2 were plated onto glass bottom dishes (WillCo) coated with Matrigel ( in Nutristem XF/FF medium (Biological Industries) supplemented for 24 hours with 10 µM ROCK inhibitor (Y-27632, Sigma-Aldrich). After 2 days, medium was replaced with KSR medium (DMEM/F12 ; 15% knockout serum replacement; 1X Glutamax; 1X non-essential amino acids; 1X penicillin-streptomycin; all from Life Technologies). This is considered day 0 of differentiation. From day 4 to day 10, KSR medium was gradually shifted to N2M medium (DMEM/F12; 1X N2 supplement; 1X Glutamax; 1X non-essential amino acids; 1X penicillin-streptomycin; all from Life Technologies). 10 µM SB431542 (Miltenyi Biotec) and 500 nM LDN193189 (Sigma-Aldrich) were added from day 0 to day 7. From day 10 onwards the medium was supplemented with 20 ng/ml BDNF, 20 ng/ml GDNF (PeproTech) and 0.2 mM ascorbic acid were added from day 10 onwards.
@@ IV.2 Immunofluorescence staining
IV.3 Time-lapse and time-course Microscopy
IV.3.1 Microscopy Setup.
We used two microscopy setups: a spinning disk microscope and a Custom Timelapse Epifluorescence Microscope. Rosettes emerge sparsely distributed at random locations on the culture plate. For this reason, we selected two microscopy setup that allowed us to follow large areas of the culture plate. Both microscope were equipped with stage top incubator. The spinning disk setup had a motorized stage and allowed the acquisition of multiple ROIs and large mosaic images. The custom microscopy was optimized for long 11-days time lapse acquisitions. Spinning Disk. This setup is an Olympus microscope with a 20x objective equipped with a stage-top CO2 incubator designed for acquiring time-lapse sequences over a period of 1-2 weeks and a motorized stage. This setup allowed us to perform time-course experiments on two culture plates simultaneously, and time-lapse experiments on a single plate. Custom Timelapse Epifluorescence Microscope. This setup is composed of an in-house assembled costume epifluorescence microscope with a 10x objective and a stage-top CO2 incubator designed for acquiring time-lapse sequences over 1-2 weeks. This set-up results in a 2 mm field width.
IV.3.2 Timecourse
The time-course experiments were performed on the spinning disk microscope. The time-course experiment allowed us to acquire large mosaics from 2 different culture plates, 3 times a day. Monitoring a large area of the culture plate, we captured the emergence of more than one rosette at a time.
IV.3.3 Timelapse
We had two different time-lapse protocols, one for the custom microscope and one for the spinning disk microscope. The time-lapse experiment used for the evaluation of the dynamical cell features was performed at the spinning disk microscope. Pictures were taken every 10 min, where in a cycle of 5 acquisitions, 1 acquisition was a bright field picture, and the other 4 were epifluorescence acquisitions on the RFP channel. This experiment was also used for the estimation of the relation between the cell displacement and the direction of the main axis of elongation by using the BF and RFP channel images.
@@ IV.4 Single cell tracking
Cell nuclei were tracked using a frame-to-frame approach. First, we used the nuclear fluorescence obtained with the reporter RFP protein expressed on the FUS protein to detect single cell nuclei. Second, the nuclei detected in each frame were linked to the nuclei in the successive frames with links using a nuclei-nuclei distance cost function. Frame-to-frame tracking was implemented using the linear assignment problem (LAP) method [22,23].
IV.5 Diffusion Analysis
From these trajectories, we obtained the different time intervals ($\Delta t$) and displacements ($\Delta x$) pairs, which we used to estimate the persistence time and length. Trajectories obtained with the tracking pipeline described above were used to extract pairs of time intervals ($\Delta t$) and displacements ($\Delta x$), enabling estimation of dynamical quantities such as the Brownian exponent $\alpha$, the motility coefficient $v_{0}$ and the correlation time $\tau$. Coordinates were first converted from pixels to microns ($1,\mathrm{pixel}=0.650,\mu\mathrm{m}$).
Firstly, considering the behaviour of each cell’s positions $(x_{t},y_{y})$ versus the time, data were plotted on a log-log scale by displaying cell displacements $\Delta r$ against time intervals $\Delta t$, revealing two distinct regimes.
Starting from the initial one, the system behaved as ballistic, for which the mean squared displacement (MSD) scales quadratically with time: $\langle\Delta r^{2}(t)\rangle\;\approx\;v_{0}^{2}\,t^{2}$ where $v_{0}$ is the self-propulsion speed or motility and $\alpha=2$. The latter one was diffusive regime, for which the MSD scales linearly with time: $\langle\Delta r^{2}(t)\rangle\;\approx\;v_{0}\,t$.
The early-time motility coefficient $v_{0}$ was quantified by fitting short-time data to the power-law model $(\Delta r)^{2}\approx v_{0}^{\alpha}\,(\Delta t)^{\alpha}$. Natural logarithms were taken, $2\ln(\Delta r)=\alpha\bigl{[}\ln(v_{0})+\ln(\Delta t)\bigr{]}$ and division by 2 yielded $\ln(\Delta r)=\beta_{1}\,\ln(\Delta t)+\beta_{0}$ where $\beta_{1}=\alpha/2$ and $\beta_{0}=(\alpha/2),\ln(v_{0})$. A linear least-squares regression was then applied and the exponent and motility coefficient were obtained from $\alpha=2\beta_{1}$ and $v_{0}=\exp\bigl{(}2\beta_{0}/\alpha\bigr{)}=\exp\bigl{(}\beta_{0}/\beta_{1}\bigr{)}$. The correlation time $\tau$, defined as the point of intersection between ballistic and diffusive power laws, was calculated as:$\tau=\exp\Bigl{(}\frac{\ln(v_{0,2})-\ln(v_{0,1})}{\alpha_{1}-\alpha_{2}}\Bigr{)}$, where $(\alpha_{1},v_{0,1})$ and $(\alpha_{2},v_{0,2})$ were derived from separate log-log regressions on data segments partitioned at a crossover time $t_{c}$ chosen to minimize combined mean-squared error. For each trajectory, values of $v_{0,i}$ and $\tau_{i}$ were compiled across the cell population, and bar plots of means with standard errors were generated to visualize their distributions.
IV.6 Elongation Analysis
[…]
IV.7 Voronoi Tesselation and structure factor analysis
@@
Voronoi tessellations were generated from fluorescence microscopy images of nuclei by manually selecting each nucleus’s center and computing the corresponding Voronoi diagram using custom Python code built on the SciPy library. Regions of interest were manually segmented and saved as label images to isolate neural‐rosette formation within each field of view, and the coordinates of the selected nuclei centers within that ROI were loaded into the Voronoi function to produce a set of polygonal regions. A zero‐valued matrix matching the input image dimensions was initialized to store the target shape factor, defined as the ratio of a region’s perimeter to the square root of its area $S_{t}=\frac{p_{i}}{\sqrt{A_{i}}}$, at the appropriate pixel locations. Infinite Voronoi regions were discarded automatically, and only polygons entirely contained within the image boundaries were retained for analysis.
To investigate the temporal evolution of cell morphology under identical culture conditions, we performed two parallel experiments—one in which neural rosette formation was observed and one in which it failed to occur. Fluorescence images were acquired daily from approximately day 6 through day 11 and cropped to a consistent 400 × 400-pixel ROI; however, significant differences between the two conditions only became apparent from day 8 onward. At each time point, nuclei centroids were identified and used to compute the target shape factor for each associated Voronoi region.
IV.7.1 Rosette identification
@@
Rosettes were manually identified analyzing the acquired mosaics. Note that while rosette emergence conditions are being understand, it is difficult to predict where they will actually form. Thus, the detection of rosette emergence in 2D cell coltures required the acquisition of large image mosaics that balance spatial resolution with extended field of view. This results in large mosaic images that must be manually analysed and hard to process and visualize. Future work should be done to address this technical issue using multiresolution data formats and visualization tools, e.g. the Cesium ion 3D tiles via ATON framework Fanini et al. (2021) which is part of the the H2IOSC infrastructure.
IV.8 Simulation details
To perform the dynamics, we followed the same procedure used for MD simulations of molecules/particle systems Grassmann et al. (2024): we first generate a system with a nucleus in random positions inside a squared box of edge $L=\sqrt{N}$ with periodic boundary conditions (to obtain them with replicate the system in every direction). Next, each simulation has been carried on for a total time of $10^{4}$, integrating equations of motions with a time step of $10^{-3}$. Forces are evaluated as done in Miotto et al. Miotto et al. (2024).

### v2

@@ III Discussions
In this work, we identified two key ingredients driving the formation of rosettes: local curvature and self-alignment interactions. The role of curvature can be understood by analyzing the configurations associated with the standard Vertex energy in Eq. 3 for a population of cells in a box with periodic boundary conditions, i.e., on a toroid.
In such a flat Euclidean space and with homogeneous parameters, the energetically optimal tiling for the cells is the hexagonal one. In fact, the number of vertices ($V$), edges ($E$), and cells ($N$) are linked by the Euler’s rule as: $V-E+N=\chi$, where $\chi$ is the Euler characteristic Del Bono et al. (2024), equal to 0 in flat topologies.
If there are no peculiar heterogeneous distributions of line tensions among the cells, each vertex connects three edges, as T-junctions are optimal to balance the edges’ tensions symmetrically. It follows that $3V=2E$, and it is straightforward to prove that the average number of edges per cell must be 6, leading to a preference for hexagonal patterns.
Regular hexagonal tiling actually corresponds to a zero Vertex energy for a value of the target shape index of $s_{0}\simeq 3.72$. Different values of shape index introduce topological frustrations in the system, which are compensated by the formation of cells with pentagonal or heptagonal shapes (see polygon distributions in Figure 1). Each deviation from hexagonal geometry inserts a topological charge that may be compensated via the formation of >3>3-fold vertices. In fact, their presence reduces the average number of sides per face. Cell activity is a promising means to modulate frustration in the cell layer. In particular, by tuning cell self-propulsion velocity, the system is known to display a solid to fluid-like behavior Bi et al. (2016). However, such ingredients are not sufficient to account for the observed presence of rosettes (see Figure 1), as in flat surfaces, even four-edge junctions are unstable configurations and rapidly result in T1 transitions, without special conditions, i.e., peculiar disposition of the line tensions.
@@
Indeed, from Plateau’s laws for soap films, we know that the angles at a vertex depend on (i) the number of edges meeting, (ii) the tensions in each edge, and (iii) the local curvature of the space. Thus, higher-fold junctions could, in principle, become favorable for specific anisotropic tension configurations, but also where curvature affects the local angular sum at vertices. Specifically, a positive curved space corresponds to a sum of angles exceeding $360^{\circ}$. As the natural meeting angles between edges change, in this scenario, rosettes become stable because distributing larger angles over more edges reduces mechanical stress. Moreover, positive curvature induces cells to tighten, which in turn produces cells with a low number of sides (see polygon distributions in Figure 1 and  2).
@@
Introducing the external harmonic field via Eq. 4, we inserted the simplest positive correction to the curvature. To accommodate such curvature, the system tends to introduce topological defects that locally reduce angular excess, forming rosette-like structures (see snapshots in Figure 2). Euler’s characteristic becomes $\chi>0$, modifying the balance of defects: an excess of positive topological charge is required.
@@
In ideal, infinite systems with flat geometry, translational symmetry ensures that the structure can be shifted in space without altering its physical properties. On curved surfaces, there is no global translational symmetry, as each point has different geometric properties. The introduction of a local curvature and the accompanying defects break this symmetry. Defects such as 5- and 7-fold polygons are therefore indicators of the breaking of translational invariance.
@@
In the present work, we produced such a breaking of the space translational invariance via either a potential energy landscape or a leader cell with a different size with respect to the rest of the population. In the former case, the external potentials used can mimic the collective effects of apical actin filament networks. Indeed, actin filaments play a fundamental role in shaping the mechanical environment of cells, particularly by forming networks that exert tensile forces across the apical surface of epithelial cell layers. These supra-cellular actin networks can generate coordinated mechanical tension, contributing to tissue morphogenesis and the maintenance of epithelial integrity Fernandez-Gonzalez and Zallen (2011). In particular, it has been demonstrated how apical actomyosin cables and medial actin structures contribute to mechanical force distribution across cell layers Rauzi et al. (2010); Fernandez-Gonzalez and Zallen (2011). Similarly, the potential can simulate the effect of an actual curved surface where cells are constrained to move, like in embryos and tumor spheroids Trichas et al. (2012). Note that in these real systems, there is a spontaneous breaking of the symmetry as internal cell mechanisms are responsible for processes like the accumulation of actin fibers or the shape of the embryo. Indeed, we show that the presence of a large cell in the layer suffices to insert a topological defect that provokes the formation of transient rosettes.
@@
Relying on a dynamical model, we were able to probe the stabilizing/destabilizing effects of cellular properties on the formation of rosettes. In particular, we found that cell motility and deformability influence the formation and maturation of rosettes. Our simulations highlighted an interplay between topological curvature, self-propulsion velocity, and target shape factor (Figure 2b-c). In the presence of fast-moving cells, rosette formation needs both an increase in space curvature and a higher cell deformability. Similarly, both the size and polarization of lumen-centered structures are favored when the cell population is in a fluid-like state. Notably, analyzing the behavior of populations of iPS-derived neural cells growing on a 2D layer via extensive fluorescent microscopy time lapses, we confirmed our model predictions, as rosettes form in regions characterized by slow-moving cells having a shape factor higher than 3.9, i.e., being in a fluid-like regime. Looking at the motion of isolated cells, we observed that their motion is maximum when it aligns with the direction of their elongation (Figure 6). Self-alignment interactions that produce cell displacements in the same direction along which cells are elongated are the second ingredient we identified that increases the probability of transient rosette formation (see Figure 3) and the size of lumen-centered ones (see Figure 4). In particular, it allows cells to orderly dispose in multiple layers around the lumen as shown in Figure 5.
@@
In conclusion, our work proved that rosettes are natural outcomes of geometric and mechanical responses to curvature, providing an additional framework where the interplay between defects and spatial curvature offers profound insights into how materials and surfaces are structured, with deep implications across disciplines as diverse as condensed matter physics and cosmology. To test our prediction, we measured the dynamical features of populations of iPSCs and found that neural rosettes are observed when cells populate the region of the parameter space our model predicts to be compatible with their formation.
@@ IV.1 Cell Culture
Human iPSCs from the WT I line Lenzi et al. (2015) were stably transfected with a construct encoding an RFP-tagged version of the nuclear FUS protein De Santis et al. (2019). iPSCs were induced to differentiate along the neuroectodermal lineage to form neural rosettes as follows. 10.000 cells/cm2 were plated onto glass-bottom dishes (WillCo) coated with Matrigel ( in Nutristem XF/FF medium (Biological Industries) supplemented for 24 hours with 10 µM ROCK inhibitor (Y-27632, Sigma-Aldrich). After 2 days, medium was replaced with KSR medium (DMEM/F12 ; 15% knockout serum replacement; 1X Glutamax; 1X non-essential amino acids; 1X penicillin-streptomycin; all from Life Technologies). This is considered day 0 of differentiation. From day 4 to day 10, KSR medium was gradually shifted to N2M medium (DMEM/F12; 1X N2 supplement; 1X Glutamax; 1X non-essential amino acids; 1X penicillin-streptomycin; all from Life Technologies). 10 µM SB431542 (Miltenyi Biotec) and 500 nM LDN193189 (Sigma-Aldrich) were added from day 0 to day 7. From day 10 onwards the medium was supplemented with 20 ng/ml BDNF, 20 ng/ml GDNF (PeproTech) and 0.2 mM ascorbic acid were added from day 10 onwards.
@@ IV.2 Immunofluorescence staining
IV.3 Microscopy experiments
Measurements were performed via an Olympus microscope with a 20x objective equipped with a V1 Spinning Disk (CrestOptics) and a stage-top CO2 incubator. Time-lapse experiments. Time-lapse experiments were used for the evaluation of the dynamical features of iPSCs when plated at low density. Pictures were taken every 10 min, where in a cycle of 5 acquisitions, 1 acquisition was a bright field picture (BF), and the other 4 were fluorescence acquisitions on the RFP channel (RFP). Time-course experiments. Time-course experiments were used to follow the formation of rosettes. Since rosettes emerge sparsely distributed at random locations on the culture plate, we acquired large mosaics from 2 different culture plates. Specifically, fluorescence images were acquired daily from day 6 through day 11 of the differentiation process. Monitoring a large area of the culture plate, we captured the emergence of more than one rosette at a time.
@@ IV.4 Single cell tracking
To measure the dynamical features of a single cell from the time-lapses, nuclei were tracked using a frame-to-frame approach. First, we used the nuclear fluorescence obtained with the reporter RFP protein expressed on the FUS protein to detect single cell nuclei. Second, the nuclei detected in each frame were connected to the nuclei in the successive frames with links using a nuclei-nuclei distance cost function. Frame-to-frame tracking was implemented using the linear assignment problem (LAP) method.
Diffusion Analysis. Trajectories obtained with the tracking pipeline described above were used to evaluate the mean square displacements. Ballistic and diffusive regimes were identified as the best fit regions for $\langle\Delta r^{2}(t)\rangle\;\approx\;v_{0}^{2}\,t^{2}$ and $\langle\Delta r^{2}(t)\rangle\;\approx\;A\,t$ where $v_{0}$ is the self-propulsion velocity. The persistence time $\tau$ has been obtained as the point of intersection between ballistic and diffusive regimes.
Elongation Analysis. To compute the elongation axis of the cells, a directional intensity analysis was implemented on BF images via a custom Python pipeline. Nuclear fluorescence was used to determine the centroid of each nucleus, from which the elongation axis was inferred. Evaluation was performed along a discrete set of orientations $\theta\in[0,\pi)$, uniformly spaced. For each direction, a linear segment of $\pm 150$ pixels (300 pixels in total) was sampled symmetrically from the centroid. The average BF intensity was computed along each segment. In BF imaging, cell boundaries typically produce high-intensity (bright) signals due to light scattering at refractive index gradients from the medium to the cell. Conversely, internal regions and elongated protrusions often present lower contrast and appear darker; the underlying assumption is that directions intersecting prominent cell boundaries will yield higher mean intensity values, whereas directions aligned with elongation, where the cell extends more continuously, will intersect fewer such transitions, resulting in lower average intensities. Thus, the direction along which the average intensity is minimized is taken as the morphological axis of maximal extension. The method provides an axial orientation and is particularly suitable for detecting elongated geometries. To complement the intensity‐based orientation, a purely geometric analysis was performed on each cell. Nuclei centroids were used to seed a watershed segmentation on the corresponding BF images, yielding one binary mask per cell. Only masks that remained fully contained within the image frame at all time points were retained to avoid boundary artifacts. For each mask, we performed a Principal Component Analysis on the covariance of the mask’s pixel positions. The first principal component was used to define the major axis of the ellipse that best fits the cell shape. From the tracking pipeline, the direction of the instantaneous displacement, $\alpha_{t}$ is obtained. The geometric elongation‐axis angle, $\beta$, was measured as in the same reference. Finally, the orientation alignment was defined by the difference $|\alpha_{t}-\beta|$.
IV.5 Voronoi Tesselation and structure factor analysis
@@
Time-course fluorescence images were cropped to a consistent 400 × 400-pixel ROI for all acquired time points. Voronoi tessellations were generated from fluorescence microscopy images of nuclei by manually selecting each nucleus’s center and computing the corresponding Voronoi diagram using custom Python code built on the SciPy library. Regions of interest were manually segmented and saved as label images to isolate neural‐rosette formation within each field of view, and the coordinates of the selected nuclei centers were used to compute Voronoi tesselation. The average target shape factor has been computed as the ratio of a region’s perimeter to the square root of its area $S_{t}=\frac{p_{i}}{\sqrt{A_{i}}}$.Infinite Voronoi regions were discarded automatically, and only polygons entirely contained within the image boundaries were retained for analysis.
@@
Rosette identification. Rosettes were manually identified by analyzing the acquired mosaics. Note that while rosette emergence conditions are being understood, it is difficult to predict where they will actually form. Thus, the detection of rosette emergence in 2D cell cultures required the acquisition of large image mosaics that balance spatial resolution with an extended field of view. This results in large mosaic images that must be manually analysed and are hard to process and visualize.
IV.6 Simulation details
We performed Molecular Dynamics simulations of the self-propelled Voronoi model by integrating numerically the stochastic equation for each cell center $\mathbf{r}_{i}$ using the Euler-Maruyama scheme. As an initial condition, we consider a random uniform distribution of $N$ cell centers arranged in a square box of side $L=\sqrt{N}$ with periodic boundary conditions. Next, each simulation has been carried out considering $N=400cells$, for a total time of $10^{4}$. Equations of motion have been integrated with a time step of $dt=10^{-3}$. Other used dynamical parameters are specified when corresponding results are discussed. We monitored that the system reached a stationary state by looking at the behavior of the descriptors used to describe rosettes ($P_{r}$, $l_{1}$, etc). Typically, stationary configurations are reached after $\sim 10^{4}$ steps. The algorithm for the analytical computation of forces within the Voronoi model is the one adopted in  Miotto et al. (2024), which is the standard fashion for force computation within the Voronoi model of biological tissues Bi et al. (2016); Sussman (2017); Barton et al. (2017).